OpenStudy (anonymous):

An express an local train leave Grayslake at 3pm and head for chicago 50 miles away. the express travels twice as fast as the local and arrives 1 hour ahead of the local. Find the speed of each train.

OpenStudy (anonymous):

I can't figure out the equation, once i have that i can solve the problem

OpenStudy (anonymous):

To tackle this questions let us first define speed. Speed is d/t, or distance over time. We can call the speed of the express train V1. So, V1 = 50/t. We can call the speed of the local train V2. So, V2 = 50/(t+1), because it take 1 additional hour to travel the same 50 miles. So, we have defined two equations but we have three unknowns. We will need one additional equation to solve. That is provided by comparing the speed of the trains. We are told V1=2*V2. Now, with substitution and the three equations we can solve.